As easy as this one seems, I have no idea where to start. Your assistance would be greatly appreciated:
A bone-in, 12 lb. ham contains 42 luncheon servings. What is the rate in servings per pound of ham?

Answer by Alan3354(21608)

A bone-in, 12 lb. ham contains 42 luncheon servings. What is the rate in servings per pound of ham?
--------------------
bone-in or bone-out has nothing to do with the problem. Sometimes things are added to cause confusion. The problem could have stated what day it was, but that's also irrelevant.
--------------
There are 42 servings from 12 pounds.
The trick is to use the units, 42 servings/12 pounds
That's 42/12 servings per pound, = 42/12 = 7/2 = 3.5
3.5 servings per pound.